OpenStudy (anonymous):

A person leaves home at 8:00 a.m. and drives to a destination at a rate of 40 mph. The person returns at a rate of 25 mph and arrives at 2:30 p.m. If x represents how far it is to the destination, then which of the following expressions represents the amount of time to make the return trip?

hero (hero):

It makes sense to include the possible expressions since there is more than one way of doing this.

OpenStudy (anonymous):

okay...one sec.

OpenStudy (anonymous):

a) 25x b) 6.5-x c) x/25

hero (hero):

For some reason I'm getting \(t = \frac{x}{25} - 6.5\)

hero (hero):

Which is none of those

hero (hero):

Chero, is there a typo?

OpenStudy (anonymous):

let me check...

OpenStudy (anonymous):

No typos on the possible answers.

OpenStudy (anonymous):

\[Rate = \frac{Distance}{Time}\] \[Time*Rate = Time*\frac{Distance}{Time}\] \[Time*Rate = Distance\] \[\frac{Time*Rate}{Rate} = \frac{Distance}{Rate}\] \[Time = \frac{Distance}{Rate}\]

OpenStudy (anonymous):

I copied and pasted the question.

OpenStudy (anonymous):

Substitute Distance and Rate to get the Time.

hero (hero):

Wired, do you have an actual explanation or are you going to continue to talk in general?

OpenStudy (anonymous):

That IS the explanation. Replace Distance and Rate and you have the answer.

hero (hero):

I see. I was focused too much on setting up the problem to solve for t

OpenStudy (anonymous):

so is the answer b??

OpenStudy (anonymous):

I am terrible at word problems....they give me headaches.

OpenStudy (anonymous):

What is the Distance? What is the Rate for the trip home?

OpenStudy (anonymous):

Since the total time: 12:30 - 8= 6.5 hrs -> Return time : total - leaving time = ....

OpenStudy (anonymous):

x/25 should be the answer. I am not sure I just went ahead and submitted my answers.

OpenStudy (anonymous):

That's the correct answer.

hero (hero):

Yes, Chero, no one wanted to give you the answer though

OpenStudy (anonymous):

I am too tired, taking a quick break!! I appreciate your help. I am sure I will be back in about 20 minutes!

OpenStudy (anonymous):

@Chero Take a look at my explanation!

OpenStudy (anonymous):

Thank you again! I have been helping my 15 year old son since noon today and am about fried!! LOL

hero (hero):

Chrorophyll, they wanted the expression for time

OpenStudy (anonymous):

That what it is: Return time = total time - start time

OpenStudy (anonymous):

Starting time = x

OpenStudy (anonymous):

@Chlorophyll Starting time is NOT X. As per the original statement: X represents how far it is to the destination.

OpenStudy (anonymous):

"If x represents how far it is to the destination"

OpenStudy (anonymous):

Yes, I just stated that. X is DISTANCE, not TIME.

OpenStudy (anonymous):

Oh, I misread it :")

OpenStudy (anonymous):